# How Developers Interact with AI: A Taxonomy of Human-AI Collaboration in Software Engineering

### v1

@@ Abstract
Artificial intelligence (AI), including large language models and generative AI, is emerging as a significant force in software development, offering developers powerful tools that span the entire development lifecycle. Although software engineering research has extensively studied AI tools in software development, the specific types of interactions between developers and these AI-powered tools have only recently begun to receive attention. Understanding and improving these interactions has the potential to improve productivity, trust, and efficiency in AI-driven workflows. In this paper, we propose a taxonomy of interaction types between developers and AI tools, identifying eleven distinct interaction types, such as auto-complete code suggestions, command-driven actions, and conversational assistance. Building on this taxonomy, we outline a research agenda focused on optimizing AI interactions, improving developer control, and addressing trust and usability challenges in AI-assisted development. By establishing a structured foundation for studying developer-AI interactions, this paper aims to stimulate research on creating more effective, adaptive AI tools for software development.
@@ I Introduction
Artificial intelligence (AI) is rapidly transforming software development, bringing new capabilities and efficiency to every stage of the development lifecycle [1]. As AI tools become increasingly sophisticated, they offer developers a diverse range of support options, from autocompletion and code generation to documentation, testing, and project management [2]. The study of developer interactions with traditional software tools has a long and rich history in software engineering research [3], providing insights into how tool usability, integration, and feedback loops shape productivity and code quality. However, the ways in which AI-powered tools can effectively support developer workflows have only recently begun to receive focused research attention [4].
This emerging focus on AI-powered tools brings with it an opportunity to explore how the interaction of the tools can be tailored to better align with developers’ needs, ultimately enhancing adoption, trust, and productivity [5]. The diversity of interaction types – from auto-complete code suggestions to conversational exchanges [6] – introduces significant complexity in designing and evaluating these tools. Without a structured framework for analyzing these interactions, researchers and tool designers lack a common vocabulary and conceptual model for discussing, comparing, and improving developer-AI interfaces. This fragmentation hinders efforts to optimize how developers engage with AI tools, identify best practices, understand usage patterns, and address common challenges across different tools and contexts. A comprehensive taxonomy of developer-AI interactions can help bridge this gap by providing a foundation for empirical studies, enabling systematic tool evaluation, and guiding the development of AI assistance features. Furthermore, as AI capabilities continue to evolve rapidly, such a taxonomy becomes essential for tracking how interaction patterns adapt and emerge over time, ensuring that new tools and features align with developers’ actual needs and workflows.
In this paper, we propose a taxonomy of interaction types between human developers and AI tools. To the best of our knowledge, this is the first taxonomy to specifically examine human-AI interaction types within the context of software engineering, a domain characterized by the need to manage diverse types of artifacts (e.g., code, bug reports, pull requests), navigate complex human collaboration dynamics, and address rapidly evolving development workflows. Building on this taxonomy, we outline a research agenda highlighting areas for further exploration, including optimizing interaction styles, improving developer control, and addressing trust and usability in AI-driven development. This paper aims to catalyze further research and discussion on creating more effective and adaptive AI tools for software development.
@@ II Interaction Types
Table I provides an overview of the proposed 11 types of developer-AI interactions, characterized by their triggers, the AI response, the developer reaction, the type of output generated, and concrete examples. The trigger column describes how each interaction is initiated, ranging from automatic mechanisms such as typing context to explicit commands or workflow events. The AI response column details the nature of the system’s output, such as real-time suggestions, context-specific recommendations, or step-by-step explanations. The developer response column captures how developers interact with and react to AI output, including actions such as reviewing, refining, or integrating suggestions. The output column highlights the type of deliverables produced by AI, such as code snippets, documentation, or quality reports, offering insight into the specific contributions these interactions make to the development process. Finally, the example column grounds each interaction type with a real-world application, illustrating the use of AI in tools like GitHub Copilot, ChatGPT, or GitLab Auto DevOps. In the following, we briefly introduce each interaction type.
@@
Conversational assistance enables developers to engage in natural language interactions with AI, making it an approachable collaborator. Developers can ask questions or request guidance on specific challenges, and the AI responds with explanations, suggestions, or even detailed code snippets. This interactive exchange facilitates learning, problem solving, and creativity, as developers adapt and refine AI responses to fit their needs. By mimicking human-like communication, this type of interaction reduces the cognitive burden of translating ideas into formal syntax, allowing developers to focus more on their core problem-solving objectives rather than wrestling with technical documentation or rigid query formats.
@@
Contextual recommendations leverage the AI’s ability to interpret cues from the project environment, providing developers with tailored suggestions aligned with the specific context of their work. For example, the AI might suggest libraries, patterns, or improvements based on the file type, project dependencies, or overall structure. Unlike auto-complete code suggestions, which are triggered in real-time as the developer types and are generally focused on immediate, localized code completion, contextual recommendations take a broader view, offering guidance that aligns with the entire project or file. By focusing on project-wide context rather than line-by-line interactions, contextual recommendations complement autocomplete suggestions to provide a more comprehensive level of support. Developers evaluate and integrate these suggestions as needed, ensuring that they align with the specific requirements of the task at hand.
@@
Comment-guided prompts allow developers to influence AI behavior using natural language descriptions embedded directly in the code as comments. The AI interprets these prompts to generate code snippets that align with the described intent. Developers can then review, adjust, and verify the generated output to ensure it meets their requirements.
@@
Event-based triggers enable automation during key workflow milestones, such as commits or pull requests. The AI responds to these events by performing predefined tasks like running quality checks or security scans and producing actionable outputs such as reports or flagged issues. Developers then review these outputs to address any identified concerns. This interaction type is used for example to integrate AI into continuous integration and delivery pipelines, automating repetitive tasks and ensuring code quality.
@@
Shortcut-activated commands streamline the interaction process by allowing developers to invoke AI functionality with predefined keyboard shortcuts. The AI responds with contextually relevant suggestions or documentation that developers can evaluate and incorporate into their work. This approach reduces disruption, allowing developers to maintain their workflow while accessing AI support quickly and efficiently.
@@
File-aware suggestions provide developers with context-specific recommendations tailored to the type of file or directory they are working on. For example, the AI might suggest configuration options for a settings file or test templates for a testing script. Developers can review these suggestions and adapt them to fit their project requirements, ensuring that AI assistance is relevant and actionable.
@@
Automated API responses integrate AI capabilities into broader project workflows through API calls or webhooks. These interactions typically involve the generation of reports, release notes, or analysis outputs in response to predefined triggers. Developers review, modify, and integrate these outputs as needed, making this interaction type particularly useful for managing large-scale or repetitive tasks within collaborative projects.
@@ IV Conclusion
This paper presents a taxonomy of the types of developer-AI interaction, providing a structured framework for understanding the diverse ways developers engage with AI-powered tools in software engineering. The accompanying research agenda builds on this taxonomy, identifying critical areas such as trust, cognitive load, and customization that warrant further exploration. Our work provides a foundation for understanding the diverse ways developers interact with AI-powered tools in software engineering. The taxonomy outlines key interaction types and dimensions, which can help guide further research and tool development. Future work could include a systematic evaluation of existing tools to analyze the frequency and usage of each interaction type. In addition, researchers could explore how these interaction types align with developer needs in specific contexts, refining the taxonomy based on empirical findings.

### v2

@@ Abstract
Artificial intelligence (AI), including large language models and generative AI, is emerging as a significant force in software development, offering developers powerful tools that span the entire development lifecycle. Although software engineering research has extensively studied AI tools in software development, the specific types of interactions between developers and these AI-powered tools have only recently begun to receive attention. Understanding and improving these interactions has the potential to enhance productivity, trust, and efficiency in AI-driven workflows. In this paper, we propose a taxonomy of interaction types between developers and AI tools, identifying eleven distinct interaction types, such as auto-complete code suggestions, command-driven actions, and conversational assistance. Building on this taxonomy, we outline a research agenda focused on optimizing AI interactions, improving developer control, and addressing trust and usability challenges in AI-assisted development. By establishing a structured foundation for studying developer-AI interactions, this paper aims to stimulate research on creating more effective, adaptive AI tools for software development.
@@ I Introduction
Artificial intelligence (AI) is rapidly transforming software development, bringing new capabilities and efficiency to every stage of the development lifecycle [1]. As AI tools become increasingly sophisticated, they offer developers a diverse range of support options, from autocompletion and code generation to documentation, testing, and project management [2]. Related work on developer interactions with traditional tools has established the importance of usability and integration [3], but there is limited research specifically on AI-powered tools[4]. The paper aims to address this gap.
This emerging focus on AI-powered tools brings with it an opportunity to explore how the interaction of the tools can be tailored to better align with developers’ needs, ultimately enhancing adoption, trust, and productivity [5]. The diversity of interaction types – from auto-complete code suggestions to conversational exchanges [6] – introduces significant complexity in designing and evaluating these tools. Without a structured framework for analyzing these interactions, researchers and tool designers lack a common vocabulary and conceptual model for developer-AI interfaces. This fragmentation hinders efforts to optimize how developers engage with AI tools, identify best practices, understand usage patterns, and address common challenges across different tools and contexts. A comprehensive taxonomy of developer-AI interactions can help bridge this gap by providing a foundation for empirical studies, enabling systematic tool evaluation, and guiding the development of AI assistance features. Furthermore, as AI capabilities continue to evolve rapidly, such a taxonomy becomes essential for tracking how interaction patterns adapt and emerge over time, ensuring that new tools and features align with developers’ actual needs and workflows.
In this paper, we propose a taxonomy of interaction types between human developers and AI tools. To the best of our knowledge, this is the first taxonomy to specifically examine human-AI interaction types within the context of software engineering, a domain characterized by the need to manage diverse types of artifacts (e.g., code, bug reports, pull requests), navigate complex human collaboration dynamics, and address rapidly evolving development workflows. Building on this taxonomy, we outline a research agenda that highlights areas for further exploration, including optimizing interaction styles, improving developer control, and addressing trust and usability in AI-driven development. This paper aims to catalyze further research and discussion on creating more effective and adaptive AI tools for software development.
@@ II Interaction Types
Table I provides an overview of the 11 proposed types of developer-AI interactions, characterized by their triggers, the AI response, the developer reaction, the type of output generated, and concrete examples. The trigger column describes how each interaction is initiated, ranging from automatic mechanisms such as typing context to explicit commands or workflow events. The AI response column details the nature of the system’s output, such as real-time suggestions, context-specific recommendations, or step-by-step explanations. The developer response column captures how developers interact with and react to AI output, including actions such as reviewing, refining, or integrating suggestions. The output column highlights the type of deliverables produced by AI, such as code snippets, documentation, or quality reports, offering insight into the specific contributions these interactions make to the development process. The example column illustrates each interaction type with real-world applications, such as GitHub Copilot and ChatGPT. In the following, we briefly introduce each interaction type.
@@
Conversational assistance enables developers to engage in natural language interactions with AI, making it an approachable collaborator. Developers can ask questions or request guidance on specific challenges, and the AI responds with explanations, suggestions, or even detailed code snippets. This interactive exchange facilitates learning, problem solving, and creativity. By mimicking human-like communication, this type of interaction reduces the cognitive burden of translating ideas into formal syntax, allowing developers to focus more on their core problem-solving objectives rather than wrestling with technical documentation or rigid query formats.
@@
Contextual recommendations leverage the AI’s ability to interpret cues from the project environment, providing developers with tailored suggestions that align with the specific context of their work. For example, the AI might suggest libraries, patterns, or improvements based on the file type, project dependencies, or overall structure. Unlike auto-complete code suggestions, which are triggered in real-time as the developer types and are focused on immediate, localized code completion, contextual recommendations take a broader view, offering guidance that aligns with the entire project or file. By focusing on project-wide context rather than line-by-line interactions, contextual recommendations complement autocomplete suggestions to provide more comprehensive support. Developers evaluate and integrate these suggestions as needed, ensuring that they align with the specific requirements of the task at hand.
@@
Comment-guided prompts allow developers to influence AI behavior using natural language descriptions embedded directly in the code as comments. The AI interprets these prompts to generate code snippets that align with the described intent. Developers can then review, adjust, and verify the generated output to ensure that it meets their requirements.
@@
Event-based triggers enable automation during key workflow milestones, such as commits or pull requests. The AI responds to these events by performing predefined tasks such as running quality checks or security scans and producing actionable outputs such as reports or flagged issues. Developers then review these outputs to address any identified concerns. This interaction type is used, for example, to integrate AI into continuous integration and delivery pipelines, automating repetitive tasks and ensuring code quality.
@@
Shortcut-activated commands streamline the interaction process by allowing developers to invoke AI functionality with predefined keyboard shortcuts. The AI responds with contextually relevant suggestions or documentation that developers can evaluate and incorporate into their work. This approach reduces disruption, allowing developers to maintain their workflow while accessing AI support quickly and efficiently. In contrast, explicit UI actions require more deliberate engagement through IDE buttons or menus.
@@
File-aware suggestions provide developers with context-specific recommendations tailored to the type of file or directory they are working on. For example, the AI might suggest configuration options for a settings file or test templates for a testing script. Developers can review these suggestions and adapt them to fit their project requirements, ensuring that AI assistance is relevant and actionable. Note that file-aware suggestions provide recommendations tailored to specific file types at the time of file creation or editing. In contrast, contextual recommendations focus on broader project-level cues and suggest improvements during ongoing development.
@@
Automated API responses integrate AI capabilities into broader project workflows through API calls or webhooks. These interactions typically involve the generation of reports, release notes, or analysis outputs in response to predefined triggers. Developers review, modify, and integrate these outputs as needed, making this interaction type particularly useful for managing large-scale or repetitive tasks within collaborative projects. Automated API responses focus on system-to-system interactions triggered by webhooks or API calls, whereas event-based triggers are specific to workflow events like commits or pull requests within the developer’s environment.
@@ IV Conclusion
This paper presents a taxonomy of the types of developer-AI interaction, providing a structured framework for understanding the diverse ways developers engage with AI-powered tools in software engineering. The accompanying research agenda builds on this taxonomy, identifying critical areas such as trust, cognitive load, and customization that warrant further exploration. Our work provides a foundation for understanding the diverse ways developers interact with AI-powered tools in software engineering. The taxonomy outlines key interaction types and dimensions, which can help guide further research and tool development, e.g., to better orchestrate various interaction types throughout the development workflow. Although validation is beyond the scope of this paper, future studies could empirically evaluate the taxonomy by observing tool use in real-world software development workflows. Future work could also include a systematic evaluation of existing tools to analyze the frequency and usage of each interaction type. In addition, researchers could explore how these interaction types align with developer needs in specific contexts, refining the taxonomy based on empirical findings.